# Affordance-Guided Reinforcement Learning via Visual Prompting

### v3

@@ IV-C Real-World fine-tuning with Dense Rewards
In our real-world experiments, we evaluate RoboFuME and KAGI after fine-tuning for K online environment training steps, with rewards computed in pixel image space. The results are shown in Table I. For each task, RoboFuME fine-tuning improves over the offline RL policies by -%, indicating further online fine-tuning is beneficial. Notably, we see lower success rates on Cube Stacking, which requires high precision and is considerably harder than the pick-place tasks in RoboFuME [24]. With KAGI, the performance further increases across the board.
While the increase in success rates achieved by KAGI over RoboFuME is modest, we notice qualitatively different behaviors learned by each policy. As an example, Fig. 6 shows a representative trajectory demonstrating policy performance for Spatula Pick-Place. The policy fine-tuned with RoboFuME, while fairly successful, commonly demonstrated the behavior where the spatula was dropped onto the plate from a height, rather than lowering and placing the spatula like the demonstrations. The robot gripper also continues moving rightward rather than staying above the placement point, indicating some coincidental successes. In comparison, KAGI’s policy fine-tuned with dense shaping rewards demonstrated behavior that matched the demonstrations more closely. For more examples comparing qualitative behavior across different tasks, please see our supplementary video.
@@ IV-D Robustness to Reduced In-Domain Demonstrations
Cloth Covering | 45% | 45% | 70% |
Spatula Pick-Place | 10% | 15% | 40% |
Cube Stacking | 5% | 15% | 25% |
Impact of reduced demos on the policy. Unsurprisingly, the offline RL policy performs worse than when trained on the standard quantity of in-domain demonstrations. Fine-tuning for K steps using RoboFuME marginally improves success rate over the offline RL policies, but we see a significant gap compared to its original performance with the standard number of demonstrations. For example, compared to its original success rate of % on Cloth Covering with demos, RoboFuME only succeeds % of the time when given demos. Therefore, the performance of RoboFuME depends quite heavily on the amount of in-domain demonstrations provided. However, KAGI reaches its original performance even with this reduction. On the Cloth Covering task, which KAGI succeeded on % of the time with demos, KAGI achieves a % success rate with demos. This small drop indicates that, even with fewer in-domain demonstrations, KAGI can recover comparable performance with more fine-tuning. Because KAGI can tolerate fewer task-specific demonstrations, it can be applied to new tasks more easily, with a smaller data burden.
Impact of reduced demos on the success classifier. The reduction in in-domain data also impacts the behavior of the VLM-based (MiniGPT-4) success classifier used by both RoboFuME and KAGI. We observed the sparse reward predictions of MiniGPT-4 to be significantly worse with less in-domain data to fine-tune the success predictor on. To mitigate this issue and eliminate the confounding factor of inaccurate sparse rewards, we modify RoboFuME’s [24] original sparse reward computation such that the reward predictor took four task-completion prompts as input (e.g., ‘Is the spatula on the plate?’, ‘Has the spatula been moved to the plate?’, etc.), and must reach a consensus across all prompts to generate a sparse reward of . This approach reduces the number of false positives, which RL algorithms often exploit. This reveals another source of fragility in the system, as fine-tuning MiniGPT-4 is heavily reliant on in-domain demonstrations to generate accurate sparse rewards crucial for online RL. Leveraging larger models like GPT-4V that have stronger zero-shot performance to extract sparse task completion rewards may help circumvent this issue.
@@
Cloth Covering | 90% | 50% | 75% |
Spatula Pick-Place | 70% | 30% | 45% |
Cube Stacking | 60% | 15% | 30% |
@@ IV-E Importance of Online Fine-tuning
Since the VLM prompting component of KAGI is inspired by MOKA [32], we additionally evaluate MOKA to understand whether online fine-tuning offers any advantages. MOKA [32] leverages GPT-4V to predict affordance keypoints and plan a sequence of action primitives (e.g., ‘lift’, ‘reach grasp’, ‘grasp’), then directly executes the action primitive sequence by computing the actions required to reach each point. We evaluate two versions of MOKA: the first with precise resets that closely match the inputs to the VLM, and the second with slight perturbations to these initial conditions that is representative of environment resets after backward policy rollouts. The purpose of this comparison is to understand whether the VLM-generated keypoints can be reused for different initial conditions in both methods, since the cost of querying GPT-4V for every new condition is prohibitively high. Compared to this, KAGI only queries for keypoints and waypoint trajectories once at the start of online training. In Table III, we see that MOKA with precise resets succeeds between to of the time depending on the task, which actually is better than the performance of our policy. However, without these precise resets, the performance of MOKA drops noticeably, which means the GPT-4V keypoints are not reusable for this method when there are even slight perturbations to the environment, since actions are computed directly with respect to these keypoints. However, KAGI can recover some of this performance loss with online fine-tuning, without the cost of additional GPT-4V queries. KAGI leverages the benefits of closed-loop systems that are more generalizable and robust to environment changes, as opposed to open-loop systems that use hand-designed action primitives. This makes, KAGI more suitable for the autonomous RL paradigm, which involves potentially imperfect policy rollouts on a real robot.
@@ V Conclusion & Future Work
We demonstrate that dense shaping rewards extracted from VLMs can help speed up online RL, and facilitate generalization to new tasks where relying only on sparse rewards is less efficient. Leveraging both dense and sparse rewards improves autonomous learning, with better robustness to reduced in-domain demonstrations than sparse rewards alone. Our reward formulation is a possible modification to make existing fine-tuning methods more robust and data-efficient. Our work has demonstrated the benefits of VLM-based dense shaping rewards, and opens up new avenues of exploration to harness the generalization capabilities of VLMs to enhance the robustness of robot learning systems.
@@
There are several areas for future work. First, sparse task completion rewards only track if the object reached the target location or not; incorporating waypoint trajectories for object tracking into the dense reward computation could further expedite learning. In addition, running large models like GPT-4V after every episode to generate rewards might be more accurate, but is extremely costly and suffers from latency issues, and future work can explore this tradeoff. Finally, further hyperparameter tuning in the dense reward computation and CalQL algorithm could improve the success rate of policies fine-tuned with dense and sparse rewards. Overall, these research directions can help efficiently scale autonomous RL systems to a greater diversity of complex tasks with different objects and environments.

### v4

@@ IV-C Real-World Fine-Tuning with Dense Rewards
In our real-world experiments, we evaluate RoboFuME and KAGI after fine-tuning for K online environment training steps, with rewards computed in pixel image space. The results are shown in Table I. For each task, RoboFuME fine-tuning improves over the offline RL policies by -%, which is consistent with the performance in [6] and confirms that further online fine-tuning is beneficial. Notably, we see lower success rates on Cube Stacking, which requires high precision and is considerably harder than all the pick-place tasks in RoboFuME [6]. With KAGI, task performance is similar or further increases across the board.
While the increase in success rates achieved by KAGI over RoboFuME is modest using the standard number of in-domain demosntrations for pre-training, we notice qualitatively different behaviors learned by each policy. As an example, Fig. 6 shows a representative trajectory demonstrating policy performance for Spatula Pick-Place. The policy fine-tuned with RoboFuME, while fairly successful, commonly demonstrated the behavior of dropping the spatula onto the plate from a height, rather than lowering and placing the spatula like the demonstrations. The robot gripper also continues moving rightward rather than staying above the placement point, indicating some coincidental successes. In comparison, KAGI’s policy fine-tuned with dense shaping rewards showed behavior that matched the demonstrations more closely. For more examples comparing qualitative behavior across tasks, please see our supplementary video.
@@ IV-D Robustness to Reduced In-Domain Demonstrations
| Cloth Covering | 45% | 50% | 75% |
| Almond Sweeping | 40% | 55% | 80% |
| Spatula Pick-Place | 10% | 30% | 60% |
| Cube Stacking | 5% | 15% | 40% |
Impact of reduced demos on the policy. Unsurprisingly, the offline RL policy performs worse than when trained on the standard quantity of in-domain demonstrations. Fine-tuning for K steps using RoboFuME somewhat improves success rate over offline RL policies, but we see a significant gap with less demos compared to its original performance with the standard number of demonstrations. Therefore, the performance of RoboFuME depends heavily on the amount of in-domain demonstrations provided. However, across all tasks, KAGI reaches close to its original performance even with this reduction. This small drop indicates that, even with fewer in-domain demonstrations, KAGI can recover comparable performance with more fine-tuning, as it uses dense shaping rewards to guide its behavior. Since KAGI can tolerate fewer task-specific demonstrations, it can be applied to new tasks more easily, with a smaller data burden.
Impact of reduced demos on the success classifier. The reduction in in-domain data crucially impacts the behavior of the VLM-based (MiniGPT-4) success classifier used by both RoboFuME and KAGI. We observed the sparse reward predictions of MiniGPT-4 were significantly worse, indicating fine-tuning the success predictor is reliant on substantial high-quality in-domain data to generate accurate sparse rewards. To mitigate this issue and eliminate the confounding factor of inaccurate sparse rewards, we modify RoboFuME’s [6] original sparse reward computation: the reward predictor was provided four task-completion prompts as input (e.g., ‘Is the spatula on the plate?’, ‘Has the spatula been moved to the plate?’, etc.), and must reach a consensus across all prompts to generate a sparse reward of . This reduces the number of false positives, which RL algorithms often exploit. The decreased accuracy of the sparse reward predictor is a key factor behind the minimal improvements of RoboFuME over offline RL, revealing another source of fragility in the system and underscoring the importance of dense shaping rewards for data-efficient fine-tuning.
@@
| Cloth Covering | 90% | 50% | 80% |
| Almond Sweeping | 100% | 65% | 80% |
| Spatula Pick-Place | 70% | 30% | 65% |
| Cube Stacking | 60% | 15% | 45% |
@@ IV-E Importance of Online Fine-tuning
Since the VLM prompting component of KAGI is inspired by MOKA [26], we additionally evaluate MOKA to understand whether online fine-tuning offers any advantages. MOKA [26] leverages GPT-4V to predict affordance keypoints and plan a sequence of action primitives (e.g., ‘lift’, ‘reach grasp’, ‘grasp’), then directly executes the action primitive sequence by computing the actions required to reach each point. We evaluate two versions of MOKA: the first with precise resets that closely match the inputs to the VLM, and the second with slight perturbations to these initial conditions that is representative of environment resets after backward policy rollouts. The comparison seeks to highlight the benefits of closed-loop RL policies over open-loop primitives. In Table III, we see that MOKA with precise resets succeeds between to of the time depending on the task, which actually outperforms KAGI. However, without these precise resets, the performance of MOKA drops noticeably, as MOKA uses pre-defined motion primitives that are not robust to even slight environment perturbations, since actions are computed directly with respect to specific VLM-generated keypoints on the objects. However, KAGI can recover most of this performance loss with online fine-tuning using VLM-generated waypoint trajectories. KAGI leverages the benefits of closed-loop systems that are more generalizable, robust to environment changes, and exhibit retry behavior, as opposed to open-loop systems that use hand-designed action primitives. This makes KAGI more suitable for the autonomous RL paradigm, which involves potentially imperfect policy rollouts on a real robot.
@@ V Conclusion & Future Work
We demonstrate that dense shaping rewards extracted zero-shot from VLMs can help speed up online RL, and facilitate generalization to new tasks where relying only on sparse rewards is less efficient. Leveraging both dense and sparse rewards improves autonomous learning, with better robustness to reduced in-domain demonstrations than sparse rewards alone. Our reward formulation can modify existing fine-tuning methods to make them more robust and data-efficient. Our work demonstrates the benefits of VLM-based dense shaping rewards, and opens up new avenues of exploration to harness the generalization capabilities of VLMs to enhance the robustness of robot learning systems.
@@
There are two key areas for future work. First, sparse task completion rewards only track if the object reached the target location or not. While this approach was successful in KAGI, incorporating object-centric waypoint trajectories into dense reward computation could further expedite learning, especially for longer-horizon tasks. Future work on improving object-centric pixel tracking and occlusion handling would aid this approach. Second, further research into improving offline RL algorithms (CalQL) could improve the performance of base policies that are fine-tuned with dense and sparse rewards, further reducing the number of online fine-tuning steps. Overall, these directions can help efficiently scale autonomous RL systems to a greater diversity of complex tasks with different objects and environments.
Appendix
Verifying Reproduction of RoboFuME. Since our proposed dense shaping rewards to improve online policy fine-tuning are integrated into RoboFuME [6], we performed preliminary experiments on two tasks featured in [6] – Cloth Folding and Cloth Covering – and verified their successful reproduction. We also performed additional experiments verifying RoboFuME’s reliance on high-quality in-domain demonstrations with low multimodality for both pre-training RL policies and fine-tuning the sparse reward classifier, showing that eliminating in-domain demonstrations entirely was catastrophic for the system’s performance. Finally, we explore RoboFuME’s selection of Bridge data subsets [55, 56] for pre-training their offline RL policies for tabletop manipulation, which we use in our experiments.
Overall, our preliminary experiments revealed that there are robustness challenges for pipelines relying on a large number of high-quality in-domain demonstrations, and incorporating task-relevant offline data facilitates transfer to the desired task. These key findings motivated the methodological details of KAGI for improving the efficiency of online fine-tuning with VLM-generated dense shaping rewards.
Qualitative Analysis of VLM Performance. The performance of our online RL fine-tuning system is influenced by the accuracy of VLM keypoint and waypoint predictions. While MOKA [26] verifies that keypoint-based reasoning of VLMs are reliably accurate for tabletop manipulation and generalizable to new objects and tasks, the paper only provides a single top-down viewpoint to the VLM, since MOKA’s open-loop primitives compute actions based on pre-determined depth information specific to the evaluation setup, which requires manual engineering and tuning.
We observe that for closed-loop RL systems, providing dual-angle viewpoints are important to accurately generate waypoint trajectories in 3D space for dense rewards. We modify MOKA’s [26] metaprompt to accommodate this additional dimension of input, so as to generate 3D waypoints instead of 2D waypoints for KAGI’s dense rewards. We keep this metaprompt the same across all tasks.
To test the robustness of our prompting structure, in addition to GPT-4o, we analyze the outputs generated from our prompts when provided to GPT-4V (used by MOKA [26], larger but slower than GPT-4o) and Gemini 1.5. Qualitative analysis of each VLM’s outputs on our tasks determines that Gemini performs similarly and can replace GPT-4o in our framework. GPT-4V, while generally reliable, sometimes struggles to generate sensible depth movements using the side-angle view to facilitate successful completion of the task. Overall, KAGI is designed to flexibly incorporate existing and future VLMs of similar or better reasoning capability compared to GPT-4o, and we leave extensive qualitative comparisons and benchmarking frameworks for VLMs on 3D spatial reasoning tasks to future work.
Please see the Additional Information section of our project page for more details on these areas of analysis and other preliminary experiments.